OpenStudy (anonymous):

In a bag there are 2 black and 4 white identical rings. Two rings are drawn from the bag at random. Find the probability that only one of them is white.

OpenStudy (anonymous):

first one is black AND second one is white OR first one is white AND second one is black is one way to proceed

OpenStudy (anonymous):

\[\frac{2}{6}\times \frac{4}{5}+\frac{4}{6}\times \frac{1}{5}\] is one way there are others

OpenStudy (anonymous):

for example you could compute \[\frac{\binom{4}{1}\times \binom{2}{1}}{\binom{6}{2}}\]

OpenStudy (anonymous):

second way is probably snappier, gives \[\frac{8}{15}\] right away

OpenStudy (klimenkov):

@satellite73 May be there is a mistake in your second post? Can you please explain why does the second way give the right result?

OpenStudy (anonymous):

thank you :)